# Understanding Dataset Distillation via Spectral Filtering

### v1

@@ 3 The Proposed Method
Both low- and high-frequency features contribute to the training of neural networks [Frequency]. It is intuitive to preserve both information in the synthetic datasets. However, existing DD methods only have fixed filter functions and cannot capture the diverse frequency information of the real datasets, which motivates the design of our model.
Curriculum Frequency Matching
For the high-pass filter $f(\lambda)=(\lambda+\beta)^{-1}$, we can control its shape by adjusting the value of $\beta$. As the value of $\beta$ increases from zero, the proportion of low-frequency information gradually increases. When $\beta$ approaches infinity, the high-pass filter will degenerate into the all-pass filter. Therefore, we propose to assign different values of $\beta$ during distillation to enhance the diversity of the synthetic datasets. Specifically, we use a pre-defined scheduler to control the change of $\beta$, defined as:
$\beta_{t}=\beta*(1+\cos(\pi t/T))/2,$ (16)
where $\beta$ is a hyper-parameter to control the frequency band, $t$ indicates the current step, and $T$ is the maximum step.
Implement of FFC and FLC.
Advanced methods [SRe2L, G-VBSM, D3S] suggest matching the representations of real and synthetic datasets in each layer of the backbone network, rather than just matching the last layer. Following this principle, CFM calculates the FFC and FLC matrices based on the input data in each layer, which is a feature map $X$ with shape $n\times d\times h\times w$, indicating the number of images, channel, height, and weight, respectively. For the FFC matrix, we reshape the feature map into channel-level representations, and for the FLC matrix, we average the height and weight dimensions of $X$:
$\hat{X}=\text{reshape}(X)\in\mathbb{R}^{nhw\times d},\quad X^{\prime}=\text{avg}(X)\in\mathbb{R}^{n\times d}.$ (17)
@@
In practice, directly computing $\hat{X}^{\top}\hat{X}$ suffers from numerical overflow problems because the reshape operator significantly increases the number of samples from $n$ to $nhw$. Therefore, we use the covariance of $\hat{X}^{\top}\hat{X}$ and the average of ${X^{\prime}}^{\top}Y$ as substitutes for FFC and FLC to stabilize the training process of CFM:
$\Psi=\frac{1}{nhw}(\hat{X}-\bar{X})^{\top}(\hat{X}-\bar{X}),\quad\Phi=\frac{1}{n}{X^{\prime}}^{\top}Y,$ (18)
@@
where $\Psi$ and $\Phi$ denote the normalized FFC and FLC matrices, and $\bar{X}$ is the average of $\hat{X}$.
Exponential Moving Updating (EMU).
@@
The computation of $\Psi$ and $\Phi$ is affordable for real datasets but is much more complex for synthetic datasets. In practice, we can only calculate $\Psi_{s}$ and $\Phi_{s}$ in a mini-batch manner, which differs from the full-batch assumption in the theoretical analysis. To mitigate this gap, we adopt a moving update technology [G-VBSM, D3S] to cache the result of each batch and ultimately approximate the statistics of the real datasets:
$\Psi_{s}=\frac{1}{b}\Psi_{s}^{b}+(1-\frac{1}{b})\Psi_{s}^{b-1},\quad\Phi_{s}=\frac{1}{b}\Phi_{s}^{b}+(1-\frac{1}{b})\Phi_{s}^{b-1},$ (19)
@@
where $\Psi_{s}^{b}$ and $\Psi_{s}^{b-1}$ indicate the FFC matrices of the current $b$-th batch and the previous batch, respectively. The symbols of the FLC matrices are the same as above.
Final Loss Function.
@@
Recalling the analysis in Section 2, the filter functions $f$ of different methods vary differently, but the choice of $g$ is limited to $g(X^{\top}Y)=I$ or $X^{\top}Y$. The former corresponds to a filter-matching loss, while the latter leads to a signal-matching loss. In the context of CFM, we define the loss functions as:
$\mathcal{L}_{\text{filter}}$ $=\sum_{l=1}^{L}\left\|(\Psi^{l}+\beta_{t}I)^{-1}-(\Psi_{s}^{l}+\beta_{t}I)^{-1}\right\|,$ (20)
$\mathcal{L}_{\text{signal}}$ $=\sum_{l=1}^{L}\left\|(\Psi^{l}+\beta_{t}I)^{-1}\Phi^{l}-(\Psi_{s}^{l}+\beta_{t}I)^{-1}\Phi_{s}^{l}\right\|,$
@@
where $l$ is the layer index of the distillation network. By combining these two functions with the traditional classification loss, we get the final loss function:
@@
where $\eta=0.1$ is a hyperparameter for all datasets for simplicity. An overview of CFM is shown in Algorithm 1.
Algorithm 1 Curriculum Frequency Matching
2:Synthetic dataset $\mathcal{S}=(H_{s},Y_{s})$
3:Feed $H$ into $\phi$ for a forward pass
4:for layer index $l=1,\cdots,L$ do
5:     Calculate $\Psi^{l}$ and $\Phi^{l}$ based on Eq. 18
6:end for
7:for batch index $b=1,\cdots,\frac{m}{|B|}$ do
8:     Initialize $H_{s}^{b}$ with randomly sampled real data
9:     Update $\beta_{t}$ based on Eq. 16
10:     repeat
11:         Feed $H_{s}^{b}$ into $\phi$ for a forward pass
12:         for layer index $l=1,\cdots,L$ do
13:              Calculate $\Psi_{s}^{l}$ and $\Phi_{s}^{l}$ based on Eq. 19
15:         end for
16:         Calculate $\mathcal{L}_{\text{ce}}$ and back-propagate $\mathcal{L}$
17:         Update $H_{s}^{b}$
18:     until Reached the number of iteration $\mathcal{I}$
19:end for
@@ 4 Experiments
Table 2: Performance (%) of different DD methods in CIFAR-10/100. The best performance is highlighted in bold. Results are taken from the original papers, and -- indicates missing data.
@@
DM DC MTT KIP G-VBSM DWA CFM SRe2L G-VBDM DWA CFM
CIFAR-10 10 48.9±0.6 44.9±0.5 65.3±0.7 46.1±0.7 46.5±0.7 45.0±0.4 52.1±0.5 27.2±0.5 53.5±0.6 32.6±0.4 57.0±0.3
50 63.0±0.4 53.9±0.5 71.6±0.2 53.2±0.7 54.3±0.3 63.3±0.7 64.0±0.4 47.5±0.6 59.2±0.4 53.1±0.3 82.3±0.4
CIFAR-100 10 29.7±0.3 25.2±0.3 33.1±0.4 40.1±0.4 38.7±0.2 47.6±0.4 58.3±0.4 31.6±0.5 59.5±0.4 39.6±0.6 64.6±0.4
50 43.6±0.4 30.6±0.6 42.9±0.3 -- 45.7±0.4 59.0±0.1 67.1±0.3 49.5±0.3 65.0±0.5 60.3±0.5 71.4±0.2
[…]
Dataset Tiny-ImageNet ImageNet-1K
IPC 50 100 10 50 100 200
ResNet-18 SRe2[SRe2L] 41.4±0.4 49.7±0.3 21.3±0.6 46.8±0.2 52.8±0.3 57.0±0.4
CDA [CDA] 48.7 53.2 -- 53.5 58.0 63.3
G-VBSM [G-VBSM] 47.6±0.3 51.0±0.4 31.4±0.5 51.8±0.4 55.7±0.4 --
DWA [DWA] 52.8±0.2 56.0±0.2 37.9±0.2 55.2±0.2 59.2±0.3 --
CFM 54.6±0.2 57.6±0.1 38.3±0.3 57.3±0.2 60.1±0.2 63.6±0.2
ResNet-50 SRe2[SRe2L] 42.2±0.5 51.2±0.4 28.4±0.1 55.6±0.3 61.0±0.4 64.6±0.3
CDA [CDA] 50.6 55.0 -- 61.3 65.1 67.6
G-VBSM [G-VBSM] 48.7±0.2 52.1±0.3 35.4±0.8 58.7±0.3 62.2±0.3 --
DWA [DWA] 53.7±0.2 56.9±0.4 43.0±0.5 62.3±0.1 65.7±0.4 --
CFM 54.8±0.2 58.2±0.3 42.9±0.4 63.2±0.2 66.1±0.3 67.9±0.4
ResNet-101 SRe2[SRe2L] 42.5±0.2 51.5±0.3 30.9±0.1 60.8±0.5 62.8±0.2 65.9±0.3
CDA [CDA] 50.6 55.0 -- 61.6 65.9 68.4
G-VBSM [G-VBSM] 48.8±0.4 52.3±0.1 38.2±0.4 61.0±0.4 63.7±0.2 --
DWA [DWA] 54.7±0.3 57.4±0.3 46.9±0.4 63.3±0.7 66.7±0.2 --
CFM 55.3±0.3 59.2±0.3 43.0±0.3 63.5±0.2 66.9±0.3 68.8±0.3
Table 4: Cross-architecture performance (%) of different DD methods. $\ddagger$ indicates that the results are our evaluation, otherwise, we adopt the results from the original paper. The results of DWA are not reported due to the lack of source code and synthetic images.
@@
In this section, we compare the proposed method with several competitive DD methods on various datasets and settings.
4.1 Experimental Setup
Datasets.
We consider four datasets, ranging from small to large scale, including CIFAR-10/100 [CIFAR] (32$\times$32, 10/100 classes), Tiny-ImageNet [Tiny] (64$\times$64, 200 classes), and ImageNet-1K [ImageNet] (224$\times$224, 1000 classes).
Networks.
Following SR2[SRe2L], we use ResNet-18 [ResNet] as the distillation network for all datasets. For a fair comparison with traditional DD methods, we also use ConvNet-128 [DC] in CIFAR-10/100, as suggested by G-VBSM. We adopt various memory budgets for different datasets, including image per class (IPC)-10, 50, 100, and 200.
Baselines.
Traditional DD methods only worked on small-scale datasets, e.g., CIFAR-10/100, but do not perform well in large-scale datasets, e.g., ImageNet-1K. In this case, we select DM [DM], DC [DC], MTT [MTT], and KIP [KIP] as baselines in CIFAR-10/100. Moreover, we report the performance of some more recent DD methods, including SRe2[SRe2L], CDA [CDA], G-VBSM [G-VBSM], and DWA [DWA] in all datasets.
Evaluation.
In the evaluation phase, we adopt the Fast Knowledge Distillation (FKD) strategy suggested by SRe2L, which has been proven to be useful for large-scale datasets. For a fair comparison, we strictly control the hyperparameters of FKD to be the same as previous methods. See Appendix 2.1 for more details.
4.2 Quantitative Results
CIFAR-10/100.
Traditional DD methods perform well on datasets with fewer classes, i.e., CIFAR-10, and shallow models, i.e., ConvNet-128. The advanced DD methods are more effective when the dataset becomes larger and the network goes deeper. We can observe that CFM consistently outperforms all advanced DD methods by a large margin. Notably, ResNet-18 trained on the original datasets has an accuracy of 94.25% and 77.45% on CIFAR-10/100. The performance of existing methods is far from achieving optimal results, while CFM is significantly approaching, demonstrating its effectiveness in distilling low-resolution data.
Tiny-ImageNet & ImageNet-1k.
In the large-scale datasets, we use ResNet-{18, 50, 101} to evaluate the performance of different methods. In the Tiny-ImageNet part, we can find that CFM has an average of 3% performance improvement over DWA, which is the most recent method using advanced model augmentation strategies. This indicates that CFM is highly competitive with state-of-the-art methods. In the ImageNet-1k part, we can see that CFM consistently surpasses all baselines in IPC=50, 100, and 200, validating its superiority over the advanced DD methods. Moreover, we observe that DWA performs better than CFM in IPC=10. We suspect that this is because the synthetic data cannot effectively cover the useful frequency features under the limited memory budget, leading to the performance degeneration of CFM.
@@ 4.3 Cross-architecture Generalization
In addition to classification accuracy, cross-architecture generalization is also crucial for DD. Ideally, synthetic datasets should encode the knowledge of real datasets rather than over-fitting to a specific model architecture. Therefore, we use diverse networks, including ResNet-50/101, DenseNet-121 [densenet], RegNet-Y-8GF [regnet], and ConvNeXt-Tiny [convnext], to evaluate the generalization ability of different DD methods on the ImageNet-1k dataset with IPC=50. The distillation network is uniformly set to ResNet-18. Results are shown in Table 4, from which we can see that CFM achieves the best performance by a substantial margin across different architectures. We speculate that the high-frequency information is transferable across different architectures. Therefore, CFM encodes the architecture-invariant knowledge of the real datasets and exhibits powerful cross-architecture generalization ability.
Table 5: Ablation studies on the loss functions of CFM. Experiments are conducted on datasets with IPC=50.
@@
$\mathcal{L}_{\text{filter}}$ $\mathcal{L}_{\text{signal}}$ $\mathcal{L}_{\text{cls}}$ CIFAR-100 Tiny ImageNet
@@
(a) Tiny-ImageNet
(b) ImageNet-1k
Figure 3: Ablation studies on the choice of filters.

### v2

@@ 5 The Proposed Method
It is crucial to preserve both the consistency and diversity of the synthetic datasets. However, existing DD methods only have fixed filter functions and cannot capture the diverse frequency information of the real datasets, which motivates the design of our model. The framework is shown in Figure 4.
Implementation of FFC and FLC. We calculate the FFC and FLC matrices based on the feature map of each convolutional layer. Specifically, the feature map of the $l$-th layer is denoted as $X^{l}_{s}\in\mathbb{R}^{m\times d\times h\times w}$, indicating the number of synthetic images, channels, height, and width, respectively. For the FFC matrix, we reshape the feature map into channel-level representations, and for the FLC matrix, we average it along the height and width dimensions:
$\hat{X}^{l}_{s}=\text{reshape}(X^{l}_{s})\in\mathbb{R}^{mhw\times d},\quad\bar{X}^{l}_{s}=\text{avg}(X^{l}_{s})\in\mathbb{R}^{m\times d}.$ (15)
@@
In practice, directly computing the FFC matrices suffers from numerical overflow problems because the reshape operator significantly increases the number of samples from $m$ to mhwmhw. Therefore, we use the covariance matrix and mean representation as substitutes to stabilize the training process
$\Psi^{l}_{s}=\frac{1}{mhw}(\hat{X}^{l}_{s}-\bar{X}^{l}_{s})^{\top}(\hat{X}^{l}_{s}-\bar{X}^{l}_{s}),\quad\Phi^{l}_{s}=\frac{1}{m}{\bar{X}^{l^{\top}}_{s}}Y_{s},$ (16)
@@
where $\Psi^{l}_{s}$ and $\Phi^{l}_{s}$ denote the normalized FFC and FLC matrices of the synthetic dataset. Similarly, we can calculate $\Psi^{l}$ and $\Phi^{l}$ for the real dataset.
Figure 4: Framework of CFM. The real and synthetic data is first fed into a backbone to calculate their corresponding FFC and FLC matrices. The filter $f(\lambda)=(\lambda+\beta_{b})^{-1}$ is then applied on FFC to extract certain frequencies. CFM uses different values of $\beta_{b}$ in each batch to cover both the low-frequency and high-frequency information.
@@
Exponential Moving Updating (EMU). The computation of $\Psi^{l}$ and $\Phi^{l}$ is affordable for real datasets but is much more complex for synthetic datasets. In practice, we can only calculate $\Psi_{s}^{l}$ and $\Phi_{s}^{l}$ in a mini-batch manner, which differs from the full-batch assumption in the theoretical analysis. To mitigate this gap, we adopt a moving update technology (G-VBSM; D3S) to cache the result of each batch and ultimately approximate the statistics of the real datasets:
$\Psi_{s}^{l,b}=\frac{1}{b}\Psi_{s}^{l,b}+(1-\frac{1}{b})\Psi_{s}^{l,b-1},\quad\Phi_{s}^{l,b}=\frac{1}{b}\Phi_{s}^{l,b}+(1-\frac{1}{b})\Phi_{s}^{l,b-1},$ (17)
@@
where $\Psi_{s}^{l,b}$ and $\Psi_{s}^{l,b-1}$ indicate the FFC matrices of the current $b$-th batch and the previous batch, respectively. The symbols of the FLC matrices are the same as above.
Curriculum Frequency Matching. After calculating the FFC and FLC matrices, we need to design a filtering function to match the crucial frequency information between the real and synthetic datasets. Specifically, we consider a filter $f(\lambda)=(\lambda+\beta)^{-1}$, whose shape is controlled by the value of $\beta$. Generally, as $\beta$ decreases, the proportion of high-frequency information gradually increases, thus improving the diversity of synthetic datasets. However, high-frequency information also introduces additional noise, resulting in performance degradation, as shown in Figure 6. Therefore, we define a scheduler to dynamically adjust the value of $\beta$ in different batches to balance the consistency and diversity of the synthetic dataset
$\beta_{b}=\beta*(1+\cos(\pi b/B))/2,$ (18)
where $\beta$ is a hyper-parameter that controls the maximum frequency, and $B$ denotes the total number of batches. Notably, $f(\Psi_{s}^{l,b})=(\Psi_{s}^{l,b}+\beta_{b})^{-1}$ has different characteristics. For example, $\Psi_{s}^{l,1}$, corresponding to $\beta_{1}$, preserves consistency and $\Psi_{s}^{l,B}$, corresponding to $\beta_{B}$, emphasizes diversity.
@@
Loss Function. After defining the filtering function $f$, another consideration is the choice of $g$, which is restricted to $g(X^{\top}Y)=I$ or $X^{\top}Y$. The former corresponds to a filter-matching loss, while the latter leads to a signal-matching loss. In the context of CFM, we define the loss functions as:
$\mathcal{L}_{\text{filter}}$ $=\sum_{b=1}^{B}\sum_{l=1}^{L}\left\|(\Psi^{l}+\beta_{b}I)^{-1}-(\Psi_{s}^{l,b}+\beta_{b}I)^{-1}\right\|,$ (19)
$\mathcal{L}_{\text{signal}}$ $=\sum_{b=1}^{B}\sum_{l=1}^{L}\left\|(\Psi^{l}+\beta_{b}I)^{-1}\Phi^{l}-(\Psi_{s}^{l,b}+\beta_{b}I)^{-1}\Phi_{s}^{l}\right\|,$
@@
By combining these two functions with the basic classification loss, we get the final loss function:
@@
where $\eta=0.1$ is a hyperparameter for all datasets. See Appendix C for more details, such as the pseudo algorithm.
@@ 6 Experiments
Table 2: Performance (%) of different DD methods in CIFAR-10/100. The best performance is highlighted in bold. Results are taken from the original papers, and - indicates missing data.
@@
DM DC MTT FrePo G-VBSM DWA CFM SRe2L G-VBDM DWA CFM
CIFAR-10 10 48.9±0.6 44.9±0.5 65.3±0.7 65.5±0.4 46.5±0.7 45.0±0.4 52.1±0.5 27.2±0.5 53.5±0.6 32.6±0.4 57.0±0.3
50 63.0±0.4 53.9±0.5 71.6±0.2 71.7±0.2 54.3±0.3 63.3±0.7 64.0±0.4 47.5±0.6 59.2±0.4 53.1±0.3 82.3±0.4
CIFAR-100 10 29.7±0.3 25.2±0.3 33.1±0.4 42.5±0.2 38.7±0.2 47.6±0.4 58.3±0.4 31.6±0.5 59.5±0.4 39.6±0.6 64.6±0.4
50 43.6±0.4 30.6±0.6 42.9±0.3 44.3±0.2 45.7±0.4 59.0±0.1 67.1±0.3 49.5±0.3 65.0±0.5 60.3±0.5 71.4±0.2
Table 3: Performance (%) of different DD methods in Tiny-ImageNet and ImageNet-1k.
Dataset IPC ResNet-18 ResNet-101
SRe2L CDA G-VBSM RDED DWA CFM SRe2L RDED DWA CFM
T-ImageNet 50 41.4±0.4 48.8 47.6±0.3 58.2±0.1 52.8±0.2 58.0±0.2 42.5±0.2 41.2±0.1 54.7±0.3 60.4±0.2
100 49.7±0.3 53.2 51.0±0.4 - 56.0±0.2 59.2±0.1 51.5±0.3 - 57.4±0.3 61.1±0.2
ImageNet-1k 10 21.3±0.6 - 31.4±0.5 42.0±0.1 37.9±0.2 40.6±0.3 30.9±0.1 48.3±1.0 46.9±1.4 47.2±0.5
50 46.8±0.2 53.5 51.8±0.4 56.5±0.1 55.2±0.2 57.3±0.2 60.8±0.5 61.2±0.4 63.3±0.7 63.5±0.2
100 52.8±0.3 58.0 55.7±0.4 - 59.2±0.3 59.5±0.2 62.8±0.2 - 66.7±0.2 66.9±0.3
6.1 Experimental Setup
Datasets. We consider four datasets, including CIFAR-10/100 (CIFAR) ($\times32$, 10/100 classes), Tiny-ImageNet (Tiny) ($\times64$, 200 classes), and ImageNet-1K (ImageNet) ($\times224$, 1000 classes).
Network Architectures. We use ResNet-18 (ResNet) as the distillation network for all datasets. For a fair comparison with traditional DD methods, we also use ConvNet-128 (DC) in CIFAR-10/100, as suggested by G-VBSM (G-VBSM). We adopt various memory budgets for different datasets, including images per class (IPC)-10, 50, and 100.
Baselines. Traditional DD methods work on small-scale datasets, e.g., CIFAR-10/100, but do not perform well in the large-scale ImageNet-1K. In this case, we select DM (DM), DC (DC), MTT (MTT), and FrePo (FRePo) as baselines in CIFAR-10/100. Moreover, we report the performance of some more recent DD methods, including SRe2(SRe2L), CDA (CDA), G-VBSM (G-VBSM), RDED (RDED), and DWA (DWA).
Evaluation. In the evaluation phase, we adopt the Fast Knowledge Distillation (FKD) strategy suggested by SRe2L, which has been proven to be useful for large-scale datasets. For a fair comparison, we strictly control the hyperparameters of FKD to be the same as those of previous methods. See Appendix C.3 for more details.
6.2 Quantitative Results
CIFAR-10/100. Traditional DD methods perform well on datasets with fewer classes, i.e., CIFAR-10, and shallow models, i.e., ConvNet-128. The advanced DD methods are more effective when the dataset becomes larger and the network goes deeper. We can observe that CFM consistently outperforms all advanced DD methods by a large margin. Notably, ResNet-18 trained on the original datasets has an accuracy of 94.25% and 77.45% on CIFAR-10/100. The performance of existing methods is far from achieving optimal results, while CFM is significantly approaching, demonstrating its effectiveness in distilling low-resolution data.
Tiny-ImageNet & ImageNet-1k. In the large-scale datasets, we use ResNet-{18, 101} to evaluate the performance of different methods. In Tiny-ImageNet, CFM outperforms DWA by 3% on average and is slightly lower than RDED when IPC=50. When IPC increases to 100, CFM achieves the best performance, demonstrating its effectiveness. In the ImageNet-1k part, we can see that CFM consistently surpasses all training-based DD methods in IPC={50, 100}, validating its superiority over the advanced DD methods. Moreover, we observe that RDED performs better than CFM in IPC=10. We suspect that the synthetic data cannot effectively cover the useful frequency features under the limited memory budget, leading to the performance degeneration of CFM.
Table 4: Cross-architecture performance (%) of different DD methods. $\ddagger$ indicates that the results are our evaluation; otherwise, we adopt the results from the original paper. The results of DWA are not reported due to the lack of source code and synthetic images.
@@ 6.3 Cross-architecture Generalization
Table 5: Ablation studies on the loss functions of CFM. Experiments are conducted on three datasets with IPC=50.
@@
$\mathcal{L}_{\text{filter}}$ $\mathcal{L}_{\text{signal}}$ $\mathcal{L}_{\text{cls}}$ CIFAR Tiny ImageNet
@@
Table 6: Ablation studies on filters. Left: Tiny-ImageNet; Right: ImageNet-1K.
Table 7: Time overhead of three DD methods.
Time (s) $\mathcal{L}_{\text{cls}}$ $\mathcal{L}_{\text{BN}} +$$\mathcal{L}_{\text{dense}} +$$\mathcal{L}_{\text{filter}}+\mathcal{L}_{\text{signal}}$
SRe2L 0.16 0.26 - -
G-VBSM 0.16 - 0.35 -
CFM 0.16 - - 0.31
Table 8: Space overhead of three DD methods.
Space (MB) $\mathcal{L}_{\text{cls}}$ $\mathcal{L}_{\text{BN}} +$$\mathcal{L}_{\text{dense}} +$$\mathcal{L}_{\text{filter}}+\mathcal{L}_{\text{signal}}$
SRe2L 17186 22522 - -
G-VBSM 17186 - 23170 -
CFM 17186 - - 21450
In addition to classification accuracy, cross-architecture generalization is also crucial for DD. Ideally, synthetic datasets should encode the knowledge of real datasets rather than overfitting to a specific model architecture. Therefore, we use diverse networks, including ResNet-50/101, DenseNet-121 (densenet), RegNet-Y-8GF (regnet), and ConvNeXt-Tiny (convnext), to evaluate the generalization ability of different DD methods on the ImageNet-1k dataset with IPC=50. The distillation network is uniformly set to ResNet-18. Results are shown in Table 4, from which we can see that CFM achieves the best performance by a substantial margin across different architectures.